# Choose Your Model Size: Any Compression by a Single Gradient Descent

### v1

@@ 2.2 Any Compression via Iterative Pruning (ACIP)
We now present Any Compression via Iterative Pruning (ACIP), which consists of the following three key steps:
@@
Apply SVD to the weights $\mathbf{W}_{l}$ of all (dense) linear layers according to (6), and introduce low-rank adapters and singular value masks as tunable parameters.
@@
Choose a surrogate loss $\mathcal{L}$ and a calibration data set $\mathbf{X}$. Perform iterative pruning of singular values and simultaneous training of low-rank adapters by applying gradient-based optimization as in (8). Obtain a parameter score map $\boldsymbol{\rho}$ by using Algorithm 1.
@@
Choose any desired compression rate. Use score map $\boldsymbol{\rho}$ and low-rank adapters $\boldsymbol{\Delta}$ to materialize the compressed model.
We provide a more detailed description of these steps in the following. For a visual presentation of ACIP, we refer back to Figure 2.
@@
where $\mathbf{M}_{l}$ is a binary diagonal matrix with entries $\mathbf{m}^{(i)}_{l}\in\{0,1\}$ masking the singular values $\mathbf{s}_{l}^{(i)}$ in $\boldsymbol{\Sigma}$, and $\boldsymbol{\Delta}_{l}$ is a low-rank adapter (LoRA) (Hu et al., 2022). We find that the addition of a low-rank adapter helps to compensate for potential errors that are introduced through pruning. We initialize $\mathbf{M}_{l}$ to be the identity matrix and $\boldsymbol{\Delta}_{l}$ to zero weights. In this way, the reparametrized model is initialized to be identical to the original model up to numerical precision.
@@
The above parametrization leads to a parameter-efficient compression. Indeed, given an $m\times n$ matrix $\mathbf{W}$, the number of non-zero singular values is bounded by $\min(m,n)$, which means the number of tunable mask parameters scales linearly in the feature dimensions.
Mask parametrization
We parametrize binary masks through a thresholding operation of the form
@@
where $\mathbf{p}_{l}^{(i)}$ are scalar learnable parameters. As this operation is not differentiable, we use the straight-through estimator for backpropagation (Bengio et al., 2013; Yin et al., 2018).
@@
Our goal is to find a global score map over all singular values in reparametrized layers that will guide model compression in Step 3. Leveraging the sparsity-inducing property of $\ell_{1}$-regularization (see Section 2.1.3), we progressively shrink the mask parameters $\mathbf{p}_{l}^{(i)}$ to zero and build a score map based on the pruning order. As such, iterative scoring has two key components: (i) $\ell_{1}$-regularized iterative pruning, and (ii) the generation of the score map. We next describe each component in more detail.
Figure 3: Progressive shrinkage of exemplary mask parameters $\mathbf{p}$ in Attn-V layer $l=30$ of LLaMA-7B based on Equation 8. The starting points of shrinkage are predictive of the pruning order, a typical phenomenon in $\ell_{1}$-regularization. In ACIP, this pruning order determines the score of associated singular values $\mathbf{s}$ (compare Algorithm 1).
Iterative Pruning
@@
where $\mathcal{L}$ denotes a suitable calibration loss for the model, $\mathbf{p}=\{\mathbf{p}_{0},\dots,\mathbf{p}_{L}\}$ the set of all mask parameters, $\boldsymbol{\Delta}=\{\boldsymbol{\Delta}_{0},\dots,\boldsymbol{\Delta}_{L}\}$ the set of all low-rank adapters, and $\boldsymbol{\theta}$ the set of all remaining model parameters that are frozen during optimization. We perform optimization until a preset stopping compression rate $r_{\text{stop}}$ is reached (see Section D.2 for further discussion). Optionally, we perform post-tuning for a fixed number of steps by freezing the masks and continue optimizing the low-rank adapters.
@@
If $\lambda$ is chosen too small, the stopping rate $r_{\text{stop}}$ might never be reached. If $\lambda$ is too large, training might become unstable and the score map ambiguous. Therefore, we use a simple linear scheduler that scales $\lambda$ by a fixed factor $>$1 every $j$ optimization steps, so that pruning becomes increasingly aggressive over time.
From Iterative Pruning to Score Map
We use the optimization process of (8) to inform our score map. Based on the discussion of Section 2.1.3, we hypothesize that there is a close relationship between the order in which the parameters $\mathbf{p}_{l}^{(i)}$ are pruned and their importance for the model — the least important parameters are pruned first and so on. During experimentation, we observed a shrinkage behavior that supports this hypothesis (see Figure 3).
Algorithm 1 describes how the score map is updated after each optimization step. In plain words, the score map is built based on the pruning order. A negative number in the map indicates how many steps ago a parameter was pruned. For all parameters that have not been pruned, the score is set to the value of the corresponding parameter. We refer to Section D.7 for visual examples of ACIP score maps.
Algorithm 1 ensures that (i) the score map stores the pruning history, and (ii) it estimates future pruning based on the parameter magnitudes. Note that absolute values of the score are irrelevant for parameter ranking.
Algorithm 1 Score map generation
# params: parameters p, vectorized;
# scores: pruning scores, vectorized;
# initialized to NaN
def update_scores(scores, params):
# previously pruned parameters
score_mask = scores <= 0.
scores[not score_mask] = params[not score_mask]
scores[score_mask] -= 1.
return scores
@@
From the optimization of (8), we obtain two components: (i) the score map that informs model compression and (ii) the tuned low-rank adapters $\boldsymbol{\Delta}$. The pruned masks $\mathbf{m}$ are discarded, as they are irrelevant for compression at this stage (cf. Figure 2). As motivated above, the score map allows us to globally rank all singular vectors based on their score. We can flexibly create a model for any compression rate $r$ by pruning the singular values $\mathbf{s}$ (and corresponding singular vectors) according to their score $\rho_{l}^{(i)}$. Note that there is a monotonic, but non-linear relationship between the number of total pruned singular values $k$ and compression rate $r$ (cf. Section 2.1.2). Given a target rate $r$, we can easily find $k$ via a binary search.
Finally, for layers with determined rank $k\geq\frac{mn}{m+n}$, we avoid excess storage required by the reparametrization by simply recovering $\mathbf{W}$ from its SVD components.
3 Related work
Post-training compression methods for Foundation Models largely fall into three categories: knowledge distillation, quantization and (un-)structured pruning. For a comprehensive survey covering all of these, we refer to (Zhu et al., 2024) and the references therein.
ACIP represents a specific form of structured pruning: factorization-based compression. Conventional structured pruning (Frantar & Alistarh, 2023; Xia et al., 2024; Ashkboos et al., 2024) jointly removes groups of parameters (e.g., network layers or matrix columns/rows). At high compression rates, however, such coarse approaches often remove important weights, causing a significant performance drop that is only recoverable through additional fine-tuning.
Low-Rank Decomposition
Approximation of weight matrices through appropriate factorization aims to preserve critical information while being simple to implement. After initial exploration for smaller language models (Edalati et al., 2022; Tahaei et al., 2022), methods for LLMs primarily built on (weighted) SVD of linear layers (Ben Noach & Goldberg, 2020; Hsu et al., 2022). Large approximation errors, however, made additional fine-tuning on down-stream tasks necessary.
Follow-up work recognized that the poor approximations are caused by LLM weights being high-rank and instead turned to decompose network features which are sparse (Kaushal et al., 2023; Yu & Wu, 2023). Notably, ASVD (Yuan et al., 2024) explicitly accounts for the data distribution eliciting these activations and SVD-LLM (Wang et al., 2024) derives an analytical layer-wise correction.
Recent studies (Sharma et al., 2023; Yuan et al., 2024; Jaiswal et al., 2024) have shown rank reduction to differently affect layers in a network and proposed different heuristics for non-uniform pruning of singular values.
A common feature of these works is that they first truncate singular values to a desired size before computing an error correction. A change in compression ratio therefore prompts another round of computation. At the same time, the correction quality varies across different ratios, so that milder compression may not automatically preserve more predictive quality. In contrast, our approach creates consistent compression trade-offs from a single round of computation (optimization) and performs corrections on-the-fly.
Figure 4: Compression-performance trade-off curves generated by ACIP on C4. Compared to Figure 1, we plot over the actual model parameter sizes to enable cross-model comparison. Square marks denote the base model performance.
Rate-distortion theory
Rate-distortion theory investigates the analytical trade-off between achievable data compression rates and the error (distortion) introduced by general lossy compression algorithms (Cover & Thomas, 2006). While some recent work (Gao et al., 2019; Isik et al., 2022) investigates rate-distortion theory of machine learning models for simple architectures under rather specific assumptions, the information-theoretic limits of neural network compression are generally unknown in practically relevant settings. In this context, the family of compressed models generated by ACIP conveniently provides an empirical (upper) bound on the distortion-rate function of a large-scale model from a single optimization run.
@@ 4 Experiments
Experimental Setup
Table 1: Zero-shot evaluation of LLaMA-7B, comparing with existing SVD-based compression methods ASVD (Yuan et al., 2024) and SVD-LLM (Wang et al., 2024). $\uparrow$: larger is better; $\downarrow$: smaller is better; best results for each task and size ratio are marked in bold. The scores for ASVD and SVD-LLM are taken from Wang et al. (2024).
C4 ↓ WikiText-2 ↓ Openb. ↑ ARC_e ↑ WinoG. ↑ HellaS. ↑ ARC_c ↑ PIQA ↑ MathQA ↑ LM Eval Avg. ↑
Size Method

### v2

@@ 2.2 Any Compression via Iterative Pruning (ACIP)
Algorithmically, ACIP consists of the following three key steps, which are detailed in the sections below. For a schematic visualization of ACIP, we refer to Figure˜2.
@@
Apply SVD to the weights of all (dense) linear layers according to (6). Introduce low-rank adapters $\boldsymbol{\Delta}$ and singular value masks as tunable parameters.
@@
Choose a surrogate loss $\mathcal{L}$ and a calibration data set $\mathbf{X}$. Perform iterative pruning of singular value masks $\mathbf{p}$ and simultaneous tuning of low-rank adapters $\boldsymbol{\Delta}$ by applying $\ell_{1}$-regularized gradient-based optimization as in (8). Obtain a global parameter score map $\boldsymbol{\rho}$ by using Algorithm˜1.
@@
Choose any desired compression rate. Use the score map $\boldsymbol{\rho}$ and low-rank adapters $\boldsymbol{\Delta}$ to materialize the compressed model in real-time.
@@
where $l$ denotes the layer index, $\mathbf{M}_{l}$ is a diagonal matrix with binary entries $\mathbf{m}^{(i)}_{l}\in\{0,1\}$ masking the singular values $\mathbf{s}_{l}^{(i)}$ in $\boldsymbol{\Sigma}_{l}$, and $\boldsymbol{\Delta}_{l}$ is a low-rank adapter (LoRA) Hu et al. (2022). In all subsequent steps of ACIP, we freeze $\boldsymbol{\Sigma}_{l}$, $\mathbf{U}_{l}$, and $\mathbf{V}_{l}$.
We find that adding a low-rank adapter helps to compensate for potential errors that are introduced by pruning in Step 2 (Section˜2.2.2). We initialize $\mathbf{M}_{l}$ as the identity matrix and $\boldsymbol{\Delta}_{l}$ as zero weights. In this way, the reparametrized model remains identical to the original model up to numerical precision.
@@
The above parametrization leads to a parameter-efficient compression scheme. Indeed, given an $m\times n$ matrix $\mathbf{W}$, the number of non-zero singular values is bounded by $r=\min(m,n)$, which means the number of tunable mask parameters scales linearly in the feature dimensions.
Mask parametrization.
We parametrize the binary masks through a thresholding operation of the form
@@
where $\mathbf{p}_{l}^{(i)}$ are scalar learnable parameters. As this operation is not differentiable, we use the straight-through estimator for backpropagation Bengio et al. (2013); Yin et al. (2018).
@@
We now aim to build a global score map over all singular values in the reparametrized layers, which guide model compression subsequently in Step 3 (Section˜2.2.3). Leveraging the sparsity-inducing property of $\ell_{1}$-regularization (see Section˜2.1.3), we progressively shrink the mask parameters $\mathbf{p}_{l}^{(i)}$ to zero and derive a score map based on the pruning order. The two key algorithmic components of this “iterative scoring” strategy are presented next.
Iterative pruning.
@@
where $\mathcal{L}$ denotes a suitable calibration loss for the model, $\mathbf{p}=\{\mathbf{p}_{0},\dots,\mathbf{p}_{L}\}$ is the set of all mask parameters, $\boldsymbol{\Delta}=\{\boldsymbol{\Delta}_{0},\dots,\boldsymbol{\Delta}_{L}\}$ is the set of all low-rank adapters, and $\boldsymbol{\theta}$ is the set of all remaining model parameters that are frozen during optimization. We perform gradient-based optimization until a preset maximum compression ratio $r_{\text{stop}}$ is reached (see Section˜D.4 for further discussion). Optionally, we perform post-tuning for a fixed number of steps by freezing the masks $\mathbf{p}$ and continuing the optimization of the low-rank adapters $\boldsymbol{\Delta}$.
@@
If $\lambda$ is chosen too small, the maximum compression ratio $r_{\text{stop}}$ might never be reached. If $\lambda$ is too large, training might become unstable and the score map ambiguous. Therefore, we use a simple linear scheduler that scales $\lambda$ by a fixed factor >>1 every $j$ optimization steps, so that pruning becomes increasingly aggressive over time.
Remark 2.2 (Role of the Calibration Loss).
As pointed out in Section˜2.1.3, a feature importance ranking as produced by optimizing (8) depends on the specific loss $\mathcal{L}\bigl(\mathbf{X};\boldsymbol{\theta},\mathbf{p},\boldsymbol{\Delta}\bigr)$. Given a model to be compressed, this leaves two important design decisions to the user of ACIP, namely the choice of loss function $\mathcal{L}$ and calibration data. In the case of LLMs, it is natural to use the same loss function as in pre-training, i.e., the negative log-likelihood loss for next-token prediction. The calibration data should be representative of the task for which the compressed model(s) should perform well. This choice is therefore very use-case dependent and could range from very specific fine-tuning tasks to general-purpose prediction. In this work, we focus on the latter.
From iterative pruning to score map.
The optimization process of (8) is used to construct our score map. Based on the discussion of Section˜2.1.3, we hypothesize that there is a close relationship between the order in which the parameters $\mathbf{p}_{l}^{(i)}$ vanish and their importance for the model — the least important parameters are pruned first and so on. When conducting our experiments, we observed a shrinkage behavior that supports this hypothesis; see Figure˜3 for a specific example.
Algorithm˜1 describes how the score map is updated after each optimization step to represent feature importances. In plain words, the score map is built based on the pruning order. A negative number in the map indicates how many steps ago a parameter was pruned. For all parameters that have not been pruned, the score is set to the value of the corresponding parameter. We refer to Section˜D.9 for visual examples of ACIP score maps.
1# params: parameters p, vectorized;
2# scores: pruning scores, vectorized;
3# initialized to NaN
4
5def update_scores(scores, params):
6 # previously pruned parameters
7 score_mask = scores <= 0.
8 scores[not score_mask] = params[not score_mask]
9 scores[score_mask] -= 1.
10 return scores
Algorithm 1 The procedure for updating the score map after each optimization step. For unpruned mask parameters $\mathbf{p}$, the score is their current magnitude, which estimates the future pruning order. Once a parameter is pruned, its score becomes a negative integer value that tracks the pruning history by decrementing at each step. This dual mechanism establishes a global importance ranking used to compress the model to any target size in Step 3.
Figure 3: Progressive shrinkage of exemplary mask parameters $\mathbf{p}$ in Attn-V layer $l=30$ of LLaMA-7B based on (8). Each plotted line corresponds to the evolution of a parameter value over training time. The starting points of shrinkage are predictive of the pruning order, a typical phenomenon in $\ell_{1}$-regularization. In ACIP, this pruning order determines the score of associated singular values $\mathbf{s}$ (cf. Algorithm˜1).
The approach of Algorithm˜1 ensures that (i) the score map stores the pruning history, and (ii) it estimates future pruning based on the parameter magnitudes. Note that absolute values of the score are irrelevant for parameter ranking.
@@
From Step 2, we only retain the score map $\boldsymbol{\rho}$ and the low-rank adapters $\boldsymbol{\Delta}$. In particular, the pruned masks $\mathbf{m}_{l}^{(i)}$ are discarded, as they are irrelevant for compression at this stage (cf. Figure˜2). As motivated above, the score map allows us to globally rank all singular values based on their score. This leads to a fully independent compression stage where we can flexibly create a model of any reduced size: we prune as many singular values $\mathbf{s}_{l}^{(i)}$ (and the corresponding singular vectors) according to their scores $\boldsymbol{\rho}_{l}^{(i)}$ so that a given compression rate $r$ is achieved. Note that there is a monotonic but non-linear relationship between the total number of pruned singular values $k$ and compression rate $r$ (see Section˜2.1.2). Given a target rate $r$, we find $k$ via a binary search (in almost real-time). As this compression procedure operates directly on the reparametrized model from Step 1, it is reversible and therefore indeed allows for Any Compression; for a slightly refined version, see Section˜D.3.
Finally, to materialize a model at a fixed target rate, all pruned singular values and vectors are discarded, so that the initial SVD-reparametrization turns into an actual low-rank factorization. For layers with determined rank $k\geq\frac{mn}{m+n}$, i.e., where a factorization would not save any parameters, we avoid an inefficient storage usage by simply recovering the (dense) weight matrix from its SVD components.
Remark 2.3 (Merging Low-Rank Adapters).
Above, we have treated the truncated weight matrices $\tilde{\mathbf{W}}_{l}$ and the corresponding low-rank adapters $\boldsymbol{\Delta}_{l}$ as distinct components. It is possible to merge the two decompositions into a single one to improve model throughput. To this end, let $\boldsymbol{\Delta}_{l}=\mathbf{A}_{l}\mathbf{B}_{l}^{\top}$ be the factorization of the low-rank adapter. Then, we can express the reparametrization at layer $l$ in (6) by matrix concatenations of the form
$\mathbf{W}_{l}=\begin{bmatrix}\mathbf{U}_{l}\boldsymbol{\Sigma}_{l},\mathbf{A}_{l}\end{bmatrix}\begin{bmatrix}\mathbf{V}_{l}^{\top}\\ \mathbf{B}_{l}^{\top}\end{bmatrix}$. (9)
After specifying a compression rate, the corresponding weight matrix can be expressed as
$\tilde{\mathbf{W}}_{l}=\underbrace{\begin{bmatrix}\tilde{\mathbf{U}}_{l}\tilde{\boldsymbol{\Sigma}}_{l},\mathbf{A}_{l}\end{bmatrix}}_{=:\mathbf{M}_{2}}\underbrace{\begin{bmatrix}\tilde{\mathbf{V}}_{l}^{\top}\\ \mathbf{B}_{l}^{\top}\end{bmatrix}}_{=:\mathbf{M}_{1}},$ (10)
where $\tilde{\mathbf{U}}_{l}$, $\tilde{\boldsymbol{\Sigma}}_{l}$, $\tilde{\mathbf{V}}_{l}^{\top}$ are the truncated singular components, see (4). Note that $\tilde{\mathbf{W}}_{l}$ is never explicitly materialized as a dense matrix in memory (to ensure actual memory savings) and its forward pass is implemented as two consecutive vector-matrix multiplications with $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$.
2.3 Computational Considerations
In this section, we discuss the computational costs and memory overhead of ACIP and compare them to other common compression paradigms. As detailed in Section˜2.2, the overall process of ACIP can be divided into three distinct stages, each with different computational characteristics. A summary of the empirical runtime and memory costs for a LLaMA-7B base model is provided in Table˜A5.
The most resource-intensive stage of ACIP is scoring via iterative pruning (Step 2). It involves backpropagation through the entire model to update the singular value masks and the low-rank adapters. As such, these updates are parameter-efficient, but the overall memory consumption is still proportional to the full model size, which can be demanding. In contrast, backpropagation-free, layer-wise methods such as ASVD (Yuan et al., 2024) and SVD-LLM (Wang et al., 2024) are notably more memory-efficient, as they only need to process and store the data for a single layer at any given time. However, it is crucial to emphasize that the iterative pruning stage is compatible with Any Compression and represents a one-time, upfront computational investment. This calibration can be performed once by a model provider, amortizing the cost across all subsequent uses. Another noteworthy aspect of Step 2 is that the SVD-reparametrization initially leads to an increased model size compared to the original model (cf. Table˜A5), since the singular vector matrices $\mathbf{U}$ and $\mathbf{V}$ both need to be stored in memory.
Once the score map is generated, the compression stage (Step 3) is exceptionally efficient. As shown in Table˜A5, compressing the model to any target size by discarding singular vectors based on their global scores occurs in near real-time. This allows practitioners to dynamically select the optimal compression-performance trade-off for their specific application without incurring any significant computational delay.
@@ 3 Experiments
Experimental setup.
To demonstrate effectiveness across architectural differences in LLMs, we evaluate ACIP on a selection of popular open-weight models: LLaMA-7B/13B Touvron et al. (2023a), LLaMA-2-7B/13B Touvron et al. (2023b), LLaMA-3.1-8B Grattafiori et al. (2024), Qwen2.5-7B/14B Qwen et al. (2024), and Mistral-7B-v0.3 Jiang et al. (2023). We use a subset of C4 training data Raffel et al. (2019) for the pruning stage. Regarding evaluation tasks, we follow Wang et al. (2024) and report perplexity on validation held-outs of C4 Raffel et al. (2019) and WikiText-2 Merity et al. (2017), and we consider seven zero-shot tasks from EleutherAI LM Evaluation Harness (LM-Eval) Gao et al. (2023). More implementation details about ACIP and choices of hyperparameters can be found in Appendix˜B.
3.1 Analyzing Compression-Performance Trade-Offs
We first study compression-performance trade-offs powered by ACIP. Figures˜4 and 5 demonstrate smooth and consistent curve shapes for all considered models; analogous results for WikiText-2 and individual zero-shot LM-Eval tasks can be found in Figure˜A8. We note that a monotonic relationship between size and performance is not self-evident, e.g., see Figure˜A13 in Section˜D.6 for a trivial approach that uses magnitude-based pruning. More insights on the runtime and memory consumption of ACIP as well as inference speed of compressed models are reported in Table˜A5 in Section˜D.1.
A remarkable observation is that the oldest models, LLaMA-7B/13B, perform best perplexity-wise, while newer, more capable models like Qwen2.5-7B/14B dominate on LM-Eval as expected, especially on the lower compression levels. This apparent contradiction is likely caused by a deviation of the pre-training data distributions from C4 in the case of more recent models.
A second noteworthy outcome of Figures˜4 and 5 are the gaps between LLMs of different base model sizes in the same family. Indeed, ACIP cannot match the performance of base models of smaller size, e.g., compare the compressed Qwen2.5-7B with the original Qwen2.5-3B. This is not surprising because the corresponding smaller-size base models were obtained by pre-training or knowledge distillation Hinton et al. (2015); Busbridge et al. (2025), which are orders of magnitudes more expensive than ACIP.
Figure 4: Compression-performance trade-offs generated by ACIP on C4. Each curve was obtained by the Any Compression stage (Step 3 in Section˜2.2.3), i.e., no additional computation was required except for a perplexity evaluation. Square marks denote the base model performance.
3.2 Comparison to Existing Works
We now compare ACIP to recent works focusing on SVD-based structured pruning, namely ASVD Yuan et al. (2024), SVD-LLM Wang et al. (2024), and Dobi-SVD (without remapping) Qinsi et al. (2025). The former two approaches are backpropagation-free and perform (activation-aware) layer-wise updates instead, while Dobi-SVD proposes a differentiable truncation mechnism for singular values. Moreover, we evaluated a simple SVD magnitude pruning approach (SVD-Magn.), where we set the score map equal to the singular values of the weight matrices. This technique allows for Any Compression analogously to ACIP and therefore serves as another natural baseline; see Section˜D.6 for further study.
Table˜1 shows that ACIP consistently outperforms all baseline methods with a growing gap for higher compression levels. Note that SVD-LLM and Dobi-SVD were calibrated on WikiText-2 instead of C4, which might explain slightly better results on the former dataset for 70% and 80% size. We think that these results underpin the benefits of an end-to-end scheme: (i) a simultaneous correction, e.g., by LoRA, can drastically improve performance, and (ii) robust pruning patterns can be found without leveraging any specific features of the SVD factorization. Moreover, we note that re-computations are required to generate each row of Table˜1 for ASVD, SVD-LLM, and Dobi-SVD, whereas ACIP only needs a single run. Analogous results for ACIP applied to all other models can be found in Table˜A2.
Table 1: Any Compression under SVD reparameterization. Zero-shot evaluation of LLaMA-7B. Comparison with baselines SVD magnitude pruning (SVD-Magn.), ASVD Yuan et al. (2024), SVD-LLM Wang et al. (2024), and Dobi-SVD (without remapping) Qinsi et al. (2025). $\uparrow$: larger is better; $\downarrow$: smaller is better; best results for each task and size ratio are marked in bold. The scores for ASVD and SVD-LLM are taken from Wang et al. (2024) and the scores for Dobi-SVD from Qinsi et al. (2025).
C4 WikiText-2 Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA LM Eval
Size Method Avg. ↑